# Large circular photogalvanic effect in non-centrosymmetric magnetic Weyl semimetal CeAlSi

### v3

@@ Abstract
The recent discovery of the Weyl semimetal CeAlSi with simultaneous breaking of inversion and time-reversal symmetries has opened up new avenues for research into the interaction between light and topologically protected bands. In this work, we present a comprehensive examination of shift current and injection current responsible for the circular photogalvanic effect in CeAlSi using first-principles calculations. Our investigation identifies a significant injection current of 4 mA/V2 over a broad range in the near-infrared region of the electromagnetic spectrum, exceeding previously reported findings. In addition, we explored several externally controllable parameters to further enhance the photocurrent. A substantial boost in the injection current is observed when applying uniaxial strain along the $c$-axis of the crystal – a 5% strain results in a remarkable 64% increment. The exceptional photocurrent response in CeAlSi suggests that magnetic non-centrosymmetric Weyl semimetals may provide promising opportunities for novel photogalvanic applications.
@@ V External Influences
To account for yet another external influence, we choose strain as the parameter. We apply compressive strain on the unit cell in two difference ways: (i) along a-axis and (ii) along c-axis. In the case of (i), we observe a steep decline in the magnitude of $\eta_{xzx}$. Whereas, applying compressive stress along the c-axis drastically improves the circular photogalvanic effect. We start with the relaxed system in its ground state and gradually decrease the length of c-axis in steps of 1%. The photocurrent profiles for all those strain configurations are depicted in Fig. 7a with the legends implying the stresses corresponding to each step. The maximum value of $\eta_{xzx}$ goes up by 64% from the pristine value at 5% strain corresponding to a stress of 10.4 GPa; which implies a huge current response of 6510 $\mu$A/V2. Not only that, the plateau region of the red curve gets sharpened gradually with the application of stress and the peak gets centered in the FWHM region. This leads to a shift of the peak frequency from 0.64 eV to 0.77 eV for a stress of 10.4 GPa. The chemical poterntial in this case was kept at the Fermi level ($\epsilon_{F}$). We have also tried changing the chemical potential along with the stress, but could not observe any significant improvement. Furthermore, we have studied the shift currents under strain (Fig. 7b). However, unlike the injection current, $\sigma_{zzz}$ peaks decreased when the stresses are applied. The values away from the peaks are somewhat increased as compared to the pristine, but still remain very small relative to peak values.
@@
To understand the effect of strain on shift-current and injection current, we carefully analyze Eq. (2) and (3). The term $r^{c}_{nm;a}$ in the expression of $\sigma_{2}^{abc}$, which is defined as the generalized derivative of the Berry connection, is directly proportional to the shift vector. Shift vector is defined as $\mathcal{R}^{b}_{mn;a}=\partial_{k_{a}}\text{Arg}(A^{b}_{mn})+(A^{a}_{mm}-A^{a}_{nn})$. This vector quantifies the real-space displacement of the electrons during optical transitions [76, 77]. Localized wavefunctions cause a decrease in the shift vector, while delocalized wavefunctions do exactly the opposite. Consequently, the bunching of the wave packets under compressive strain leads to a decrease in shift current. The injection current, however, is primarily influenced by the diagonal components of the band velocities, or more precisely, the difference between the band velocities ($\Delta^{a}_{mn}$). Compressive strain increases the velocities of the electrons by making the atoms more tightly packed. There is an overall increase in $\Delta^{a}_{mn}$ for the dominant band transitions (Fig. 5). Therefore, we realize the increase in the injection current under compressive strain. Another interesting point to look at is the distribution of Weyl nodes Figs. 8(d-f). The system in its ground state holds 20 pairs of nodes with opposite chirality having energy up to $110$ meV. At 5.7 GPa stress, the nodes scatter more in energy scale and remain between $-50$ to $180$ meV. Interestingly, increasing stress beyond this point creates more Weyl points and at a stress of 10.4 GPa, there are 56 nodes in total scattered between $-100$ and $250$ meV. In addition, the degeneracies of the nodes also increase with pressure, which is indicated by the horizontal distribution of the nodes in energy landscape.

### v4

# Large circular photogalvanic effect in the noncentrosymmetric magnetic Weyl semimetal CeAlSi

@@ Abstract
The recent discovery of the Weyl semimetal CeAlSi with simultaneous breaking of inversion and time-reversal symmetries has opened up new avenues for research into the interaction between light and topologically protected bands. In this work, we present a comprehensive examination of the shift current and injection current responsible for the circular photogalvanic effect in CeAlSi using first-principles calculations. Our investigation identifies a significant injection current of 4 mA/V2 over a broad range in the near-infrared region of the electromagnetic spectrum, exceeding previously reported findings. In addition, we explored several externally controllable parameters to further enhance the photocurrent. A substantial boost in the injection current is observed when applying uniaxial strain along the $c$ axis of the crystal: a 5% strain results in a remarkable 64% increment. The exceptional photocurrent response in CeAlSi suggests that magnetic non-centrosymmetric Weyl semimetals may provide promising opportunities for novel photogalvanic applications.
DOI: 10.1103/PhysRevB.111.195105
@@ V External Influences
To account for yet another external influence, we choose strain as the parameter. We apply compressive strain on the unit cell in two difference ways: (i) along a-axis and (ii) along c-axis. In the case of (i), we observe a steep decline in the magnitude of $\eta_{xzx}$. Whereas, applying compressive stress along the c-axis drastically improves the circular photogalvanic effect. We start with the relaxed system in its ground state and gradually decrease the length of c-axis in steps of 1%. The photocurrent profiles for all those strain configurations are depicted in Fig. 7a with the legends implying the stresses corresponding to each step. The maximum value of $\eta_{xzx}$ goes up by 64% from the pristine value at 5% strain corresponding to a stress of 10.4 GPa; which implies a huge current response of 6510 $\mu$A/V2. Not only that, the plateau region of the red curve gets sharpened gradually with the application of stress and the peak gets centered in the FWHM region. This leads to a shift of the peak frequency from 0.64 eV to 0.77 eV for a stress of 10.4 GPa. It is important to note that the atomic positions were relaxed after each application of stress, and the chemical potential was kept at the Fermi level ($\epsilon_{F}$). We have also tried changing the chemical potential along with the stress, but could not observe any significant improvement. Furthermore, we have studied the shift currents under strain (Fig. 7b). However, unlike the injection current, $\sigma_{zzz}$ peaks decreased when the stresses are applied. The values away from the peaks are somewhat increased as compared to the pristine, but still remain very small relative to peak values.
@@
To understand the effect of strain on shift-current and injection current, we carefully analyze Eq. (2) and (3). The term $r^{c}_{nm;a}$ in the expression of $\sigma_{2}^{abc}$, which is defined as the generalized derivative of the Berry connection, is directly proportional to the shift vector. Shift vector is defined as $\mathcal{R}^{b}_{mn;a}=\partial_{k_{a}}\text{Arg}(A^{b}_{mn})+(A^{a}_{mm}-A^{a}_{nn})$. This vector quantifies the real-space displacement of the electrons during optical transitions [76, 77]. Localized wavefunctions cause a decrease in the shift vector, while delocalized wavefunctions do exactly the opposite. Consequently, the bunching of the wave packets under compressive strain leads to a decrease in shift current. The injection current, however, is primarily influenced by the diagonal components of the band velocities, or more precisely, the difference between the band velocities ($\Delta^{a}_{mn}$). The observed enhancement of the $xzx$-component of the injection current is due to the Poisson effect. Specifically, the uniaxial compressive strain along the $z$-direction leads to a tensile strain in the $x$-direction. This in turn results in an overall increase in $\Delta^{x}_{mn}$ for the dominant band transitions (Fig. 5). Therefore, we realize the increase in the injection current under compressive strain.Another interesting point to look at is the distribution of Weyl nodes Figs. 8(d-f). The system in its ground state holds 20 pairs of nodes with opposite chirality having energy up to $110$ meV. At 5.7 GPa stress, the nodes scatter more in energy scale and remain between $-50$ to $180$ meV. Interestingly, increasing stress beyond this point creates more Weyl points and at a stress of 10.4 GPa, there are 56 nodes in total scattered between $-100$ and $250$ meV. In addition, the degeneracies of the nodes also increase with pressure, which is indicated by the horizontal distribution of the nodes in energy landscape.